# arXiv:2603.01219v1[math.AT] 01 Mar 2026

### v1

@@ 1. Introduction
In [35] Schlessinger–Stasheff showed that the Halperin-Stasheff model leads to a representation of the set of homotopy type $(X,i:{\mathcal{H}}\cong H^{*}(X))$ as the quotient of a conical rational algebraic variety $V$ modulo a pro-unipotent group action $G$ [35]. Here $V$ is the space of perturbations of the differential in the Tate–Jozefiak resolution of $H^{*}(X)$ and $G$ is pro-unipotent algebraic group. In this formulation “finding a complete set of invariants for rational homotopy types seems to be about the same order of difficulty as finding a complete set of invariants for the $G$-orbits of a variety VV… except for special cohomology rings” [35]. Recent refinements using unital cyclic $C_{\infty}$-algebras by Kajiura [25, 26] and Hamilton–Lazarev [18] have further specialized these tools for closed oriented manifolds. Significant results include Miller’s 1979 formality result for (k1)(k-1)-connected closed manifolds of dimension less than or equal 4k24k-2 [32], using Quillen theory, and Crowley–Nordström’s 2021 classification of (r1)(r-1)-connected manifolds of dimension less than or equal (5r3)(5r-3) via the Bianchi–Massey tensor, using Sullivan theory [10].
This paper aims to refine Kadeishvili RHT classification using minimal $C_{\infty}$-algebras to define a complete set of invariants of the rational homotopy type of a topological space whose cohomology algebra is of finite type. Kadeishvili results [20], [22] stated that for a DGCA $\mathcal{A}$$^{*}$ there exists a minimal $C_{\infty}$-algebra enhancement $C_{\infty}{\mathcal{A}}^{*}$ of $H^{*}({\mathcal{A}}^{*})$ which is homotopy equivalent to $\mathcal{A}$$^{*}$. Furthermore, $C_{\infty}{\mathcal{A}}^{*}$ and $C_{\infty}{\mathcal{B}}^{*}$ are isotopic if and only if $\mathcal{A}$$^{*}$ and $\mathcal{B}$$^{*}$ are weakly equivalent.
@@
Our paper is organized as follows. In Section 2, we recall the concepts of a Hodge homotopy $d^{-}$, the corresponding harmonic subspace, and the Hodge decomposition, which is stronger than those considered by Chuang-Lazarev [7] and by Kajiura [25]. In Section 3, we show that the transferred minimal $C_{\infty}$-algebra associated with a Hodge homotopy $d^{-}$ is unital cyclic (Theorem 3.3). We also consider several symmetries of the minimal unital cyclic $C_{\infty}$-algebra, transferred using Hodge homotopy, that shall be needed in the proof of formality theorems in later sections. Section 4 is devoted to classification of real-and rational homotopy type of simply connected topological spaces, in particular, closed simply connected manifolds, using minimal $C_{\infty}$-algebra and minimal unital cyclic $C_{\infty}$-algebra enhancements and their isotopy modulo $k$, $k\in{\mathbb{N}}^{+}$, deformation theory of filtered differential graded Lie algebras (DGLAs), Harrison cohomology, and cyclic cohomology. We give explicit formulas for the first and secondary obstruction classes (Proposition 4.8, Lemmas 4.9, 4.10, Proposition 4.11). We describe the generalized affine structure of the set of complete invariants of the rational homotopy types of simply-connected topological spaces with a given cohomological algebra of finite type (Theorem 4.12, Proposition 4.16). Then we give a proof of Theorem 4.17 on the rational homotopy type of a (r1)(r-1)-connected, $r\geq 2$, manifold of dimension less than or equal l(r1)+2l(r-1)+2, $l\geq 4$. We illustrate our technique in Subsection 4.5. In Sections 5 and 6, we give proofs of formality and almost formality theorems (Theorems 5.1, 6.1), using Hodge homotopy and the resulting symmetries of the transferred minimal unital cyclic $C_{\infty}$-algebras, proved in Section 3. In the last section we discuss related results and some open questions.
@@ 2. Hodge homotopy and unital cyclic $C_{\infty}$-algebras
[13, Def. 2.1], cf. [28, Def. 4.1], cf. [11, Definition 2.1]. A PDGCA $({\mathcal{A}}^{*},d)$ of degree $n$ over a field $\mathbb{F}$} is a DGCA ${\mathcal{A}}^{*}=\oplus_{k=0}^{n}{\mathcal{A}}^{k}$ whose cohomology ring $H^{*}({\mathcal{A}})$ is finite dimensional over $\mathbb{F}$} and has a fundamental class $\int\in(H^{n}({\mathcal{A}}))^{\vee}$ such that the pairing, defined by using the cup-product,
@@
where $\cdot$] stands for the projection ${\mathcal{A}}^{n}\subseteq\ker d\to H^{n}({\mathcal{A}}^{*})$. The PDGCA $({\mathcal{A}}^{*},d)$ is called non-degenerate, [11, Definition 2.2], or oriented Poincaré duality algebra [28, Definition 2.1] if the pairing $\langle\cdot,\cdot\rangle$ on $\mathcal{A}$$^{*}$ is non-degenerate. A PDGCA $(H^{*},d=0)$ is called a Poincaré GCA.
@@
[11, Definition 2.2], [13, Definition 2.4]. The graded subspace $\mathcal{H}^{\ast}:=\mathrm{Im}(\pi_{\mathcal{H}^{\ast}})$ is called the harmonic subspace for the Hodge homotopy $d^{-}$.
@@
1) Our working example of a PDGCA over $\mathbb{R}$} admitting a Hodge homotopy is the de Rham complex $({\Omega}^{*}(M),d)$ of a closed smooth Riemann manifold $M$ with the harmonic subspace being the space of harmonic forms on $M$ [11, Remark 2.7 (1)], [13, Example 2.7]. In [11, Theorem 3.10, (4.1)] Fiorenza-Kawai–Lê–Schwachhöfer proposed a simple algorithm to find a nondegenerate finite dimensional PDGCA which is weakly equivalent to $({\Omega}^{*}(M),d)$ if $b_{1}(M)=0$.
@@ 3. Cyclicity of the minimal unital $C_{\infty}$-algebra via Hodge homotopy transfer
We prove (3.9) by induction on $k$. For $k=3$, equation (3.9) follows directly from (3.7). Assume that (3.9) holds for all indices up to (k1)(k-1). By (3.5) and by the the vanishing property (3.8) we have
@@ 4.2. Isotopy modulo $k$ for minimal $C_{\infty}$-algebra enhancements
We note that the notion of isotopy modulo $k$, $k\in{\mathbb{N}}^{+}$, on the set of minimal $C_{\infty}$-algebra enhancements can be described using a filtration on $g_{H^{*}}$ that have been considered by Hamilton-Lazarev in [18]. Namely we consider the following filtration
@@
consists of Harrison cochains of arity $\geq k$. Then we have
${\mathfrak{g}}_{H^{*}}=F^{1}{\mathfrak{g}}_{H^{*}}\supset F^{2}{\mathfrak{g}}_{H^{*}}\supset\ldots$
such that
1. $\delta(F^{p}{\mathfrak{g}}_{H^{*}})\subset F^{p}{\mathfrak{g}}_{H^{*}}$,
2. $F^{p}{\mathfrak{g}}_{H^{*}},F^{q}{\mathfrak{g}}_{H^{*}}]\subset F^{p+q}{\mathfrak{g}}_{H^{*}}$,
3. $\bigcap_{p}F^{p}{\mathfrak{g}}_{H^{*}}=0.$
Furthermore, two elements $m,m^{\prime}\in{\mathfrak{g}}_{H^{*}}$ are isotopic modulo $k$ iff there exists $p\in{\mathfrak{g}}^{0}_{H^{*}}$ such that $e^{{\rm ad}_{p}}\ast(m)=m^{\prime}\mod F^{k}{\mathfrak{g}}_{H^{*}}$. In this filtered DGLA formalism, the relation $\sim_{k}$ is effectively a group action on the quotient ${\mathfrak{g}}_{H^{*}}/F^{k}{\mathfrak{g}}_{H^{*}}$. Hence, Proposition 4.5 is immediate. Note that Hamilton-Lazarev considered minimal $C_{\infty}$-algebra structures modulo filtration and obstructions to extendability of a minimal $C_{\infty}$-algebra structure modulo $F^{n}{\mathfrak{g}}_{H^{*}}$ to a minimal $C_{\infty}$-algebra structure modulo $F^{n+1}{\mathfrak{g}}_{H^{*}}$.
@@ 2. $F^{p}{\mathfrak{g}}_{H^{*}},F^{q}{\mathfrak{g}}_{H^{*}}]\subset F^{p+q}{\mathfrak{g}}_{H^{*}}$,
4.3. The Primary Obstruction: Equivalence Modulo 3
@@ 4.4. The higher obstructions: Isotopy Modulo $k\geq 4$
Let us begin with the secondary obstruction for extending isotopy modulo 3 to isotopy modulo 4. Assume that $m\sim_{3}m^{\prime}$. Then there exists an element $p=\phi_{2}+\ldots\in F^{2}{\mathfrak{g}}^{0}_{H^{*}}$ such $e^{{\rm ad}_{p^{\prime}}}\ast m=m\mod F^{3}{\mathfrak{g}}_{H^{*}}$. By (4.16), we have
@@
(4.25) $p\in F^{2}{\mathfrak{g}}^{0}_{H^{*}},\,e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{k-1}{\mathfrak{g}}_{H^{*}}\}.$
1) Then $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ consists of cocycles in $C^{k,(2-k)}(H^{*},H^{*})$ and the image ${\mathcal{K}}_{k}(m,m^{\prime})$ of $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ in $Harr^{k,-(k-2)}(H^{*},H^{*})$ depends only on $m\mod F^{k}=(m_{3},\ldots,m_{k})$ and $m\mod F^{k}=(m_{3}^{\prime},\ldots,m_{k}^{\prime})$.
@@
$\delta\big(m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big)+\frac{1}{2}\big[m^{\prime}+\tilde{\kappa}_{k}(p,m,^{\prime}),m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\mod F^{k+1}.$
@@
$\frac{1}{2}\big[\tilde{\kappa}_{k}(p,m,m^{\prime}),\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\mod F^{k+1}.$
The first term in the above equality vanishes since $m^{\prime}$ is a MC element. The last two terms vanish since $m^{\prime}\in F^{3}$ and $\tilde{\kappa}_{k}\in F^{k}$, $k\geq 3$. It follows that $\delta\tilde{\kappa}_{k}(p,m,m^{\prime})=0\mod F^{k+1}$. This completes the proof of Lemma 4.13. ∎
@@
$(e^{\text{ad}_{q}}\ast m^{\prime})+\alpha\equiv m^{\prime}+\beta\mod{F^{k+1}}.$
@@
(4.41) $e^{\text{ad}_{q}}\ast m^{\prime}-m^{\prime}\equiv\beta-\alpha\mod{F^{k+1}}.$
@@
Let $p\in F^{2}{\mathfrak{g}}^{0}$ such that $e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{k-1}$. Set
@@
(4.53) $e^{{\rm ad}_{q}}\ast m^{\prime}-m^{\prime\prime}$ $=0\mod F^{k-1},$
@@
${\mathcal{K}}_{k}(p,m^{\prime},m^{\prime}):\big\{[\kappa_{k}(q,m^{\prime},m^{\prime})]\in Harr^{k,-(k-2)}(H^{*},H^{*}):$
$q\in F^{2}{\mathfrak{g}}_{0},\,e^{{\rm ad}_{q}}\ast m^{\prime}=m\mod F^{k}\big\}$
@@
$(e^{\text{ad}_{q}}*m^{\prime}-m^{\prime\prime})=0\mod F^{k-1}.$
@@
$\tilde{\kappa}_{k}(q*p,m,m^{\prime\prime})=\tilde{\kappa}_{k}(q,m,m^{\prime})+\tilde{\kappa}_{k}(p,m,m^{\prime})\mod F^{k}.$
@@
4.5. Example: A simply connected non-formal 8-manifold
In this subsection, we demonstrate our stratification of minimal $C_{\infty}$-algebra enhancements using the 8-dimensional simply connected symplectic manifold $\tilde{M}$ constructed by Fernández and Muñoz [12]. This manifold serves as a crucial benchmark for our theory, as it lies on the boundary of the dimension bound $n\leq 5r-2$ for $r=2$.
Recall that $\tilde{M}$ is obtained by a symplectic resolution of the orbifold $M/\mathbb{Z}_{3}$, where $M=\Gamma\backslash(H_{\mathbb{C}}\times\mathbb{C})$ is a compact nilmanifold. The Lie algebra $\mathfrak{g}$ of the complex Heisenberg group $H_{\mathbb{C}}$ has left-invariant complex 1-forms $\{\mu,\nu,\theta\}$ satisfying the structure equations:
(4.67) $d\mu=0,\quad d\nu=0,\quad d\theta=\mu\wedge\nu.$
The non-formality of the nilmanifold $M$ is well-known and is encoded in the non-triviality of the triple Massey product $\langle[\mu],[\nu],[\mu]\rangle$.

### v2

# Minimal Unital Cyclic $C_{\infty}$-Algebras and the Real and Rational Homotopy Type of Closed Manifolds

@@ Abstract.
Section 5 is written in collaboration with Domenico Fiorenza.
@@ 1. Introduction
In [35] Schlessinger–Stasheff showed that the Halperin-Stasheff model leads to a representation of the set of homotopy type $(X,i:{\mathcal{H}}\cong H^{*}(X))$ as the quotient of a conical rational algebraic variety $V$ modulo a pro-unipotent group action $G$ [35]. Here $V$ is the space of perturbations of the differential in the Tate–Jozefiak resolution of $H^{*}(X)$ and $G$ is the pro-unipotent pro-algebraic group of automorphisms of the resolution that induce the identity on cohomology. (The pro-unipotency of $G$ follows from the fact that for simply connected spaces, any automorphism inducing the identity on cohomology is the exponential of a degree-0 nilpotent derivation.) In this formulation “finding a complete set of invariants for rational homotopy types seems to be about the same order of difficulty as finding a complete set of invariants for the $G$-orbits of a variety VV… except for special cohomology rings” [35]. Recent refinements using unital cyclic $C_{\infty}$-algebras by Kajiura [25, 26] and Hamilton–Lazarev [18] have further specialized these tools for closed oriented manifolds. Significant results include Miller’s 1979 formality result for (k1)(k-1)-connected closed manifolds of dimension less than or equal 4k24k-2 [32], using Quillen theory, and Crowley–Nordström’s 2021 classification of (r1)(r-1)-connected manifolds of dimension less than or equal (5r3)(5r-3) via the Bianchi–Massey tensor, using Sullivan theory [10].
This paper aims to refine Kadeishvili RHT classification using minimal $C_{\infty}$-algebras to define a complete set of invariants of the rational homotopy type of a topological space whose cohomology algebra is of finite type. Kadeishvili results [20], [22] stated that for a DGCA $\mathcal{A}$$^{*}$ there exists a minimal $C_{\infty}$-algebra enhancement $C_{\infty}{\mathcal{A}}^{*}$ of $H^{*}({\mathcal{A}}^{*})$ which is homotopy equivalent to $\mathcal{A}$$^{*}$. Furthermore, assuming that $H^{*}({\mathcal{A}}^{*})=H^{*}({\mathcal{B}}^{*})$, then $C_{\infty}{\mathcal{A}}^{*}$ and $C_{\infty}{\mathcal{B}}^{*}$ are isotopic if and only if $\mathcal{A}$$^{*}$ and $\mathcal{B}$$^{*}$ are weakly equivalent.
@@
Our paper is organized as follows. In Section 2, we recall the concepts of a Hodge homotopy $d^{-}$, the corresponding harmonic subspace, and the Hodge decomposition, which is stronger than those considered by Chuang-Lazarev [7] and by Kajiura [25]. In Section 3, we show that the transferred minimal $C_{\infty}$-algebra associated with a Hodge homotopy $d^{-}$ is unital cyclic (Theorem 3.3). We also consider several symmetries of the minimal unital cyclic $C_{\infty}$-algebra, transferred using Hodge homotopy, that shall be needed in the proof of formality theorems in later sections. Section 4 is devoted to classification of real-and rational homotopy type of simply connected topological spaces, in particular, closed simply connected manifolds, using minimal $C_{\infty}$-algebra and minimal unital cyclic $C_{\infty}$-algebra enhancements and their isotopy modulo $k$, $k\in{\mathbb{N}}^{+}$, deformation theory of filtered differential graded Lie algebras (DGLAs), Harrison cohomology, and cyclic cohomology. We give explicit formulas for the first and secondary obstruction classes (Proposition 4.8, Lemmas 4.9, 4.10, Proposition 4.11). We describe the generalized affine structure of the set of complete invariants of the rational homotopy types of simply-connected topological spaces with a given cohomological algebra of finite type (Theorem 4.12, Proposition 4.16). Then we give a proof of Theorem 4.17 on the rational homotopy type of a (r1)(r-1)-connected, $r\geq 2$, manifold of dimension less than or equal l(r1)+2l(r-1)+2, $l\geq 4$. We illustrate our technique in Subsection 4.5. In Sections 5 and 6, we give proofs of formality and almost formality theorems (Theorems 5.1, 6.1), using Hodge homotopy and the resulting symmetries of the transferred minimal unital cyclic $C_{\infty}$-algebras, proved in Section 3. In the last section 7 we discuss related results and some open questions. This paper also contains Appendix A where we place the SageMath scripts used to verify the Fernández-Muños example.
@@
Acknowledgement
The author wishes to thank Università di Roma “La Sapienza” for hospitality and excellent working conditions during her visit in September 2025, where a part of this work was discussed. The author acknowledges the assistance of Google’s Gemini in developing the SageMath scripts used to verify the Fernández-Muñoz example in Subsection 4.5, which we place in Appendix A. Furthermore, the author thanks xAI’s Grok for symbolic calculations in verifying the trace-free Ricci cochain related to Proposition 5.6, and OpenAI’s ChatGPT for formula verification and symbolic calculations related to Proposition 6.6.
@@ 2. Hodge homotopy and unital cyclic $C_{\infty}$-algebras
[13, Def. 2.1], cf. [28, Def. 4.1], cf. [12, Definition 2.1]. A PDGCA $({\mathcal{A}}^{*},d)$ of degree $n$ over a field $\mathbb{F}$} is a DGCA ${\mathcal{A}}^{*}=\oplus_{k=0}^{n}{\mathcal{A}}^{k}$ whose cohomology ring $H^{*}({\mathcal{A}})$ is finite dimensional over $\mathbb{F}$} and has a fundamental class $\int\in(H^{n}({\mathcal{A}}))^{\vee}$ such that the pairing, defined by using the cup-product,
@@
where $\cdot$] stands for the projection ${\mathcal{A}}^{n}\subseteq\ker d\to H^{n}({\mathcal{A}}^{*})$. The PDGCA $({\mathcal{A}}^{*},d)$ is called non-degenerate, [12, Definition 2.2], or oriented Poincaré duality algebra [28, Definition 2.1] if the pairing $\langle\cdot,\cdot\rangle$ on $\mathcal{A}$$^{*}$ is non-degenerate. A PDGCA $(H^{*},d=0)$ is called a Poincaré GCA.
@@
[12, Definition 2.2], [13, Definition 2.4]. The graded subspace $\mathcal{H}^{\ast}:=\mathrm{Im}(\pi_{\mathcal{H}^{\ast}})$ is called the harmonic subspace for the Hodge homotopy $d^{-}$.
@@
1) Our working example of a PDGCA over $\mathbb{R}$} admitting a Hodge homotopy is the de Rham complex $({\Omega}^{*}(M),d)$ of a closed smooth Riemann manifold $M$ with the harmonic subspace being the space of harmonic forms on $M$ [12, Remark 2.7 (1)], [13, Example 2.7]. In [12, Theorem 3.10, (4.1)] Fiorenza-Kawai–Lê–Schwachhöfer proposed a simple algorithm to find a nondegenerate finite dimensional PDGCA which is weakly equivalent to $({\Omega}^{*}(M),d)$ if $b_{1}(M)=0$. In the case of the de Rham complex, the pairing (2.2) coincides with the integration of the wedge product of forms, and the Hodge homotopy $d^{-}$ can be realized as $Gd^{*}$, where $G$ is the Green’s operator and $d^{*}$ is the adjoint of $d$.
@@ 3. Cyclicity of the minimal unital $C_{\infty}$-algebra via Hodge homotopy transfer
We prove (3.9) by induction on $k$. For $k=3$, equation (3.9) follows directly from (3.7). Assume that (3.9) holds for all indices up to (k1)(k-1). By (3.5) and by the vanishing property (3.8) we have
@@ 4.2. Isotopy modulo $k$ for minimal $C_{\infty}$-algebra enhancements
We note that the notion of isotopy modulo $k$ ($k\in\mathbb{N}^{+}$) on the set of minimal $C_{\infty}$-algebra enhancements can be described using a filtration on the Gerstenhaber DGLA $\mathfrak{g}_{H^{*}}$ similar to those considered by Hamilton-Lazarev in [18]. Namely, we consider the arity filtration:
@@
consisting of Harrison cochains of arity $n\geq k$. This yields a descending sequence:
$\mathfrak{g}_{H^{*}}=F^{1}\mathfrak{g}_{H^{*}}\supset F^{2}\mathfrak{g}_{H^{*}}\supset\dots$
such that:
  1. (1)
    $\delta(F^{p}\mathfrak{g}_{H^{*}})\subset F^{p}\mathfrak{g}_{H^{*}}$,
  2. (2)
    $F^{p}\mathfrak{g}_{H^{*}},F^{q}\mathfrak{g}_{H^{*}}]\subset F^{p+q}\mathfrak{g}_{H^{*}}$,
  3. (3)
    $\bigcap_{p}F^{p}\mathfrak{g}_{H^{*}}=0$.
Furthermore, two elements $m,m^{\prime}\in\mathfrak{g}_{H^{*}}$ are isotopic modulo $k$ if and only if there exists a gauge element $p\in\mathfrak{g}^{0}_{H^{*}}$ such that $e^{\text{ad}_{p}}\ast m=m^{\prime}\pmod{F^{k}\mathfrak{g}_{H^{*}}}$. In this filtered DGLA formalism, the relation $\sim_{k}$ is effectively the action of the pronilpotent group associated to $\mathfrak{g}_{H^{*}}$ on the quotient $\mathfrak{g}_{H^{*}}/F^{k}\mathfrak{g}_{H^{*}}$.
Consequently, Proposition 4.5 follows immediately from the general theory of filtered DGLAs. We note that Hamilton-Lazarev utilized this framework to study obstructions to the extendibility of $C_{\infty}$-structures from $F^{n}$ to $F^{n+1}$. While the DGLA structure on Harrison cochains traces back to Gerstenhaber [14] and Voronov-Gerstenhaber [38], the emphasis on the filtered pronilpotent structure is a hallmark of the deformation-theoretic approach pioneered by Schlessinger-Stasheff [35] and Sullivan [36].
@@
4.3. The Primary Obstruction: Isotopy Modulo 3
@@ 4.4. The higher obstructions: Isotopy Modulo $k\geq 4$
Let us begin with the secondary obstruction for extending isotopy modulo 3 to isotopy modulo 4. Assume that $m\sim_{3}m^{\prime}$. Then there exists an element $p=\phi_{2}+\ldots\in F^{2}{\mathfrak{g}}^{0}_{H^{*}}$ such $e^{{\rm ad}_{p^{\prime}}}\ast m=m\pmod{F^{3}{\mathfrak{g}}_{H^{*}}}$. By (4.16), we have
@@
(4.25) $p\in F^{2}{\mathfrak{g}}^{0}_{H^{*}},\,e^{{\rm ad}_{p}}\ast m=m^{\prime}\pmod{F^{k-1}{\mathfrak{g}}_{H^{*}}}\}$.
1) Then $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ consists of cocycles in $C^{k,(2-k)}(H^{*},H^{*})$ and the image ${\mathcal{K}}_{k}(m,m^{\prime})$ of $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ in $Harr^{k,-(k-2)}(H^{*},H^{*})$ depends only on $m\pmod{F^{k}}=(m_{3},\ldots,m_{k})$ and $m\pmod{F^{k}}=(m_{3}^{\prime},\ldots,m_{k}^{\prime})$.
@@
$\delta\big(m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big)+\frac{1}{2}\big[m^{\prime}+\tilde{\kappa}_{k}(p,m,^{\prime}),m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\pmod{F^{k+1}}$.
@@
$\frac{1}{2}\big[\tilde{\kappa}_{k}(p,m,m^{\prime}),\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\pmod{F^{k+1}}$.
The first term in the above equality vanishes since $m^{\prime}$ is a MC element. The last two terms vanish since $m^{\prime}\in F^{3}$ and $\tilde{\kappa}_{k}\in F^{k}$, $k\geq 3$. It follows that $\delta\tilde{\kappa}_{k}(p,m,m^{\prime})=0\pmod{F^{k+1}}$. This completes the proof of Lemma 4.13. ∎
@@
$(e^{\text{ad}_{q}}\ast m^{\prime})+\alpha\equiv m^{\prime}+\beta\pmod{{F^{k+1}}}$.
@@
(4.41) $e^{\text{ad}_{q}}\ast m^{\prime}-m^{\prime}\equiv\beta-\alpha\pmod{{F^{k+1}}}$.
@@
Let $p\in F^{2}{\mathfrak{g}}^{0}$ such that $e^{{\rm ad}_{p}}\ast m=m^{\prime}\pmod{F^{k-1}}$. Set
@@
(4.53) $e^{{\rm ad}_{q}}\ast m^{\prime}-m^{\prime\prime}$ $=0\pmod{F^{k-1}}$,
@@
$\mathcal{K}_{k}(p,m^{\prime},m^{\prime})\coloneqq\left\{[\kappa_{k}(q,m^{\prime},m^{\prime})]\in\mathrm{Harr}^{k,-(k-2)}(H^{*},H^{*}):\begin{aligned} &q\in F^{2}\mathfrak{g}_{0},\\ &e^{\mathrm{ad}_{q}}\ast m^{\prime}=m\pmod{F^{k}}\end{aligned}\right\}$
@@
$(e^{\text{ad}_{q}}*m^{\prime}-m^{\prime\prime})=0\pmod{F^{k-1}}$.
@@
$\tilde{\kappa}_{k}(q*p,m,m^{\prime\prime})=\tilde{\kappa}_{k}(q,m,m^{\prime})+\tilde{\kappa}_{k}(p,m,m^{\prime})\pmod{F^{k}}$.
@@
4.5. Example: The Fernández-Muñoz 8-manifold